OpenStudy (anonymous):

MEDAL AND FAN!!!!! Solve the system, using substitution. What is the x coordinate of the solution? x = 2y - 2 and 3x - y = 14 A. 6 B. 4 C. 1 D. 0 E. no solution

OpenStudy (anonymous):

x = 2y - 2 3x - y = 14 ----------- 3( 2y - 2) - y = 14 6y -6 - y = 14 5y - 6 =14 5y = 20 y = 4 ----------- 3x - 4 = 14 3x = 18 x = 6